Elimination
8s+12t=16
6s+9t=12
solve

Answer by richwmiller(17219)

8s+12t=16
6s+9t=12
4*2s+4*3t=4*4
3*2s+3*3t=3*4
They are multiples of each other
There are the same line.
They share all solutions.